Guardian Inc. is trying to develop an asset-financing plan. The firm has $500,000 in temporary current assets and $400,000 in permanent current assets. Guardian also has $600,000 in fixed assets. Assume a tax rate of 30 percent.

a. Construct two alternative financing plans for Guardian. One of the plans should be conservative, with 80 percent of assets financed by long-term sources, and the other should be aggressive, with only 56.25 percent of assets financed by long-term sources. The current interest rate is 12 percent on long-term funds and 6 percent on short-term financing. Compute the annual interest payments under each plan.
Annual Interest

Conservative

Agrresive

b. Given that Guardian’s earnings before interest and taxes are $380,000, calculate earnings after taxes for each of your alternatives.
  

Earning After Taxes

Conservative

Aggresive



c. What would the annual interest and earnings after taxes for the conservative and aggressive strategies be if the short-term and long-term interest rates were reversed?
  

Conservative Agressive

Total Interest

Earning After taxes
Answer #1

(a)

Temporary current assets = $500,000

Permanent current assets = $400,000

Fixed assets = $600,000

Total assets = 500,000 + 400,000 + 600,000

= $1,500,000

Under conservative plan, 80% of the assets will be financed by long term sources.

Hence, long term funds needed = 1,500,000 x 80%

= $1,200,000

Interest on long term funds = 1,200,000 x 12%

= $144,000

Hence, remaining funds will be arranged through short term sources = 1,500,000 - 1,200,000

= $300,000

Interest on short term funds = 300,000 x 6%

= $18,000

Hence annual interest payment = 144,000 + 18,000

= $162,000

Under aggressive plan, 56.25% of the assets will be financed by long term sources.

Hence, long term funds needed = 1,500,000 x 56.25%

= $843,750

Interest on long term funds = 843,750 x 12%

= $101,250

Hence, remaining funds will be arranged through short term sources = 1,500,000 - 843,750

= $656,250

Interest on short term funds = 656,250 x 6%

= $39,375

Hence annual interest payment = 101,250 + 39,375

= $140,625

Annual interst
Conservative $162,000
Aggresive $140,625

(b)

Conservative Aggresive
Earnings before interest and tax 380,000 380,000
Less: Interest - 162,000 - 140,625
Earnings before tax 218,000 239,375
Less: Tax @ 30% - 65,400 - 71,812
Earnings after tax 152,600 167,563

(c)

Temporary current assets = $500,000

Permanent current assets = $400,000

Fixed assets = $600,000

Total assets = 500,000 + 400,000 + 600,000

= $1,500,000

Under conservative plan, 80% of the assets will be financed by long term sources.

Hence, long term funds needed = 1,500,000 x 80%

= $1,200,000

Interest on long term funds = 1,200,000 x 6%

= $72,000

Hence, remaining funds will be arranged through short term sources = 1,500,000 - 1,200,000

= $300,000

Interest on short term funds = 300,000 x 12%

= $36,000

Hence annual interest payment = 72,000 + 36,000

= $108,000

Under aggressive plan, 56.25% of the assets will be financed by long term sources.

Hence, long term funds needed = 1,500,000 x 56.25%

= $843,750

Interest on long term funds = 843,750 x 6%

= $50,625

Hence, remaining funds will be arranged through short term sources = 1,500,000 - 843,750

= $656,250

Interest on short term funds = 656,250 x 12%

= $78,750

Hence annual interest payment = 50,625 + 78,750

= $129,375

Annual interst
Conservative $108,000
Aggresive $129,375
Conservative Aggresive
Earnings before interest and tax 380,000 380,000
Less: Interest - 108,000 - 129,375
Earnings before tax 272,000 250,625
Less: Tax @ 30% - 81,600 - 75,187
Earnings after tax 190,400 175,438